OpenStudy (anonymous):

it took a small plane 2hr longer to fly 375 mi against the wind than it took the plane to fly the same distance with the wind . The rate of thewind was 25mph . Find the rate of the plane in calm air

OpenStudy (shane_b):

First, write out the system of equations based on: \[Distance = speed \space x \space time\] With the wind:\[(s+25)t=375\]Against the wind:\[(s-25)(t+2)=375\] Now you have 2 equations with 2 unknowns. You can solve this using substitution. Solving the first equation for t gets you: \[t=\frac{375}{s+25}\] Substitute that into the second equation to get: \[(s-25)(\frac{375}{s+25}+2)=375\] Solving for s you get +/- 100mph...and you know the speed is positive so it's just 100mph. You can now solve for t using the original equations: \[(100+25)t=375\]\[t = 3 hrs\] \[(100-25)(t+2)=375\]\[t = 5 hrs\]